OpenStudy (anonymous):

Find the distance between (-2, -3) and (-5, 1).

OpenStudy (jadeishere):

It's 5. this involves the Pythagoras theorem formula; a^2 + b^2 = c^2 the legs are 3 and 4 units long (a and b) so 3^2 + 4^2 = c^2 Where c is the distance you're looking for, more commonly known as the hypotenuse. so 9 + 16 = 25 the square root of 25 is 5. :) so 5 is the distance between (-2, -3) and (-5, 1)